OpenStudy (firejay5):

Name the set or sets of numbers to which each real number belongs. Use N for natural #'s, W for whole #'s, Z for integers, Q for rational #'s, and I for Irrational #'s. *Please let me know if I got any wrong and if I do explain what I did wrong/ explain the process to better aide my understanding. Also, if I forget to add one to a problem let me know as well.* 39. - (sqrt 49) 40. 0 41. 0.4583 42. 0.3 (repeating, bar notation above 3) 43. - 1/2 44. sqrt 49 45. 0.6666 46. 10/5 47. sqrt 37 48. 3.14 49. 3/5 50. 5

OpenStudy (firejay5):

My Answers: 39. Z 40. W 41. I 42. Q 43. Q 44. N, W, & Z 45. Q 46. N, W, & Z 47. I 48. Q 49. Q 50. N, W, Z, & Q

OpenStudy (jiteshmeghwal9):

For 39 ). -sqrt{49}=-7 which is both z & Q

OpenStudy (jiteshmeghwal9):

remember -7=-7/1 it is p/q form where q is not equal to 0

OpenStudy (jiteshmeghwal9):

Z={\(-\infty\),.........,-2,-1,0,1,2,.........\(\infty\)}

OpenStudy (jiteshmeghwal9):

Q=(\(-\infty \),......-2,-1,0,1,2,......\(\infty\))

OpenStudy (jiteshmeghwal9):

W=N\(\cup\){0} N={1,2,3,4,....\(\infty\)}

OpenStudy (danjs):

All of them are Rational except sqrt(37) which is irrational. All the decimals are repeating or terminate except that one

OpenStudy (firejay5):

I thought 41 was Irrational?

OpenStudy (danjs):

0.4583 It ends. you can just say... 4583/10000

OpenStudy (firejay5):

what about the even numbers? what should be the letters?

OpenStudy (danjs):

For the one Irrational, that is it, if irrational. For the rest Rational numbers, you just have to decide if Natural , whole, integer Natural -- 1,2,3,4,5... whole -- 0,1,2,3,4,5... integer-- ...-2,-1,0,1,2... Natural are counting numbers, whole adds just 0 to that, the integers add the negatives to that

OpenStudy (firejay5):

would this be the answer to the problems? 40. W & Q 42. Q 44. N, W, Z, & Q 46. N, W, Z, & Q 48. Q 50. N, W, Z, & Q

OpenStudy (danjs):

The only ones that have any of those types are 40) 0, and 50)5

OpenStudy (danjs):

So write I for 47, all the rest put a Q Then add to 40) 0 -- Z and H and add to 50) 5 --- Z , H, and N

OpenStudy (firejay5):

So I had the answers right, all I had to do was Q on all of them on the ones I didn't have a Q on them, except for 37.

OpenStudy (danjs):

Q 39. - (sqrt 49) Q,Z,H 40. 0 Q 41. 0.4583 Q 42. 0.3 (repeating, bar notation above 3) Q 43. - 1/2 Q 44. sqrt 49 Q 45. 0.6666 Q 46. 10/5 I 47. sqrt 37 Q 48. 3.14 Q 49. 3/5 Q,Z,H,N 50. 5

OpenStudy (firejay5):

H stands for what whole number's

OpenStudy (danjs):

yeah, also i forgot a couple, since those are perfect squares, and go to -7 and 7 39) Q , Z 44) Q, Z, H, N

OpenStudy (danjs):

Q,Z 39. - (sqrt 49) Q,Z,H 40. 0 Q 41. 0.4583 Q 42. 0.3 (repeating, bar notation above 3) Q 43. - 1/2 Q,Z,H,N 44. sqrt 49 Q 45. 0.6666 Q 46. 10/5 I 47. sqrt 37 Q 48. 3.14 Q 49. 3/5 Q,Z,H,N 50. 5

OpenStudy (firejay5):

oh and just making sure that 46, is N, W, Z, & Q

OpenStudy (firejay5):

@DanJS thanks for the explanation and help, it really helped me understand rational and irrational numbers on a deeper level

OpenStudy (danjs):

yeah sorry it is late , 46 is all those.. add that on

OpenStudy (firejay5):

yeah it's okay and I always appreciate a deepen explanation on how to work problems because it helps me out a lot

OpenStudy (danjs):

I would first decide Rational/irrational, then if it is an integer, natural or whole number.. Yeah, here is a nice diagram to show you. Notice how each is contained in others https://images.search.yahoo.com/yhs/search;_ylt=A0LEViOm.eBXiUoA1CknnIlQ?p=rational+and+irrational&fr=yhs-mozilla-004&fr2=piv-web&hspart=mozilla&hsimp=yhs-004#id=10&iurl=http%3A%2F%2Fleferemath.weebly.com%2Fuploads%2F1%2F4%2F1%2F9%2F14195578%2F4606013_orig.png&action=click

OpenStudy (firejay5):

omg thanks so much to be honest! :D